OpenStudy (anonymous):

Find the 20th term of the following sequence. -6, -4, -2, 0, ... 32 44 46 wouldnt this be 40??

OpenStudy (anonymous):

-6, -4, -2, 0, 2, 4, 6, 8, 10, 12, 14, 16, 18, 20, 22, 24, 26, 28, 30, 32 It would be 32

OpenStudy (ybarrap):

What's the difference between each sequence?

OpenStudy (ybarrap):

$$\tt \large \text{Then formula would be } a_n=-6+(n-1)d \text{, where d=2 and n=1,2,...}\\ \large So, \tt a_{20}=-6+(20-1)(-2)=32 $$

OpenStudy (ybarrap):

Here's a source for the formula: http://en.wikipedia.org/wiki/Arithmetic_series